OpenStudy (anonymous):

Simplified form on 3a9/18a3

OpenStudy (anonymous):

simple (1/6)a6

OpenStudy (anonymous):

\[\frac{3a^9}{18a^3}=\frac{\cancel{3}a^{9-3}}{\cancel{18}^6}\] \[=\frac{a^6}{6}\]

OpenStudy (anonymous):

my new favorite \cancel{a}